OpenStudy (nwolf):

Trains A and B are traveling int he same direction on parallel tracks. Train A is traveling at 40 mile an hour, and B is traveling at 50 miles an hour, Train A passes the station at 7:20 pm and B passes it at 7:32 pm. At what time will train B catch up to train A?

OpenStudy (anonymous):

by the drawing we have that at 7:32 both trains are separated by 8 miles (t=12 r(t)=40/60*t r(t)=4*2=8miles) now let's fix the train A as reference point. The relative displacement rate of B is Vb-Va=50-40=10miles/h or 1/6 miles/min r'(t)=1/6*t=8miles t=6*8=48minutes train B reaches A after 48 minutes

OpenStudy (anonymous):